Jennifer Kaufman bought a used Toyota Prius for $17,500. She made a $4,500 down payment and is financing the balance at a particular bank over a 3 year period at 12% interest. As her banker, calculate what equal monthly payments (in $) will be required by Jennifer to amortize the car loan. (Round your answer to the nearest cent.)
I subtracted $17,500 to $4,500 and got $13,000.
I then did 13,000 x .001/1-(1+.001)^36 = -354.83, which was wrong.

Answer by Boreal(15235)

You got a negative number, so that is the first hint something is incorrect. rate per month is 0.01 or 1%, 12%/12.
a divided by (1+r)^n-1/r(1+r)^n
n=36 payments and r is int. rate/12, for 12 payments per year.
a=$13000
the denominator is (1.01^36-1)/0.01(1.01)^36=0.431/0.143=30.107.
13000/30.107=$431.79. Note: round at the end; carry all intermediate calculations to the final one.